OpenStudy (anonymous):

Help for medal ! Rewrite the equation below in standard form 2/3x =-5/2y +2

OpenStudy (notamathgenius):

\[\frac{ 2 }{ 3 }x=- \frac{ 5 }{ 2 }y+2 \] You need to substitute. If I am correct.

OpenStudy (notamathgenius):

You need to get rid of 2/3 x so put that on both sides of the equeal sign

OpenStudy (anonymous):

Wait, this are the choices A. 4x+15y=12 B. 2x+15/2y=6 C.2x+5y=4 D. 2x+5y=6

OpenStudy (notamathgenius):



OpenStudy (notamathgenius):

 Forgot to multiply them xD